FORBIDDEN TRANSITIONS IN THE VERY RICH PURE ROTATIONAL SPECTRUM OF TRANS-1-IODOPERFLUOROPROPANE

CHRISTOPHER T. DEWBERRY, GARRY S. GRUBBS II, and \underline{STEPHEN A. COOKE}, The Department of Chemistry, University of North Texas, P.O. Box 305070, Denton, TX, 76203-5070.

Using both chirped pulse-, and low frequency cavity-based Fourier transform spectroscopy over 700 transitions have been recorded for the title molecule for the first time. The C_{1}, C_{2} and C_{3} carbon-13 species have been observed in natural abundance allowing a substitution structure for the CCC backbone to be determined. Nearly all the transitions observed were either a-type R branches or b-type Q branches. No c-type transitions were observed. The χ_{aa}, χ_{bb}, χ_{cc} and χ_{ab} components of the iodine nuclear quadrupole coupling tensor have been determined. Of note, several forbidden transitions were also observed, such as J_{k_{−1},k_{+1}} F = 14_{4,10} 23/2 ← 12_{5,7} 21/2. In this particular case it seems that the upper energy level is nearly degenerate with a level connected to the lower energy level by a μ_{a} dipole component term, but the lower energy level is also nearly degenerate with a second level connected to the upper energy level, again by a μ_{a} dipole component term. The two intermediate energy levels are themselves connected by a μ_{b} dipole component term. Near degeneracies such as these, together with a large χ_{ab} value (≈ 1 GHz) mean that certain ΔJ = 2 transitions become allowed. This phenomena has been previously observed for 1-iodopropane by Fujitake and Hayashi (J. Mol. Spect, 127, (1988), 112-124).